40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4994     Entry time: Wed Jul 20 06:17:04 2011
Author: Suresh 
Type: Update 
Category: Green Locking 
Subject: Y-end green laser power issues 

The Y-end green beam power is 0.47 mW.

While aligning the Y-end aux laser light into the fiber we noticed that the green power out of the doubling crystal was in microwatts.  I checked to see what was the trouble and found that the oven was cold as the temperature controller had been disabled.  I enabled it and scanned the temperature to maximise the green output.  Yet the power is less than 10% of that at the X end (7mW).

To verify I checked the power of various beams on the Y-end table.  They are listed below in the picture

Y-end_table_powers.pdf

The green beam power is proportional to the square of the IR incident power and this explains the drop in green power by a factor of (210/730)^2  thus making 7 mW -->  0.5 mW.  However we may be able to double the power at the Y-arm oven if the uncoated lenses in the IR path are exchaned for coated ones. 

 

The green beam injection into the Y-arm cavity also needs to be cleaned up as noted here.  As seen in the picture below two of the mirrors which launch the beam into the arm cavity need to be fixed as well.

 Y-end_table.png




ELOG V3.1.3-